What is 8y-3=15+2y
Answer by tjnw79(57)

8y-3=15+2y
first add 3 to both sides
8y-3+(3)=(3)+15+2y to get
8y=18+2y
now subtract 2y on both sides
8y-(2y)=18+2y-(2y)
6y=18
divide 6 on both sides
y=3